OpenStudy (anonymous):

f(x)=x-4 g(x)=sqrt of x+2 what is the domain of (f+g) (x) ?

OpenStudy (anonymous):

\[x \ge-2\]

OpenStudy (anonymous):

if you meant sqrt(x+2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but x≥−2 is wrong

OpenStudy (anonymous):

it's not wrong. Unless you meant sqrt(x) + 2, then the answer is x >/ 0

OpenStudy (anonymous):

got it

OpenStudy (mathstudent55):

If the functions f and g are: \(f(x)=x-4\) \( g(x)=\sqrt{x+2}\) \( (f+g)(x) = f(x) + g(x) = x - 4 + \sqrt{x + 2} \) The only restriction to the domain comes from \( \sqrt{x + 2} \) The domain is: \(x + 2 \ge 0\) \(x \ge -2\) If the functions f and g are: \(f(x)=x-4\) \( g(x)=\sqrt{x}+2\) \( (f+g)(x) = f(x) + g(x) = x - 4 + \sqrt{x} + 2 =x + \sqrt{x} - 2 \) The only restriction to the domain comes from \( \sqrt{x} \) The domain is: \(x \ge 0\)